OpenStudy (mitchi35):

Challenge: Graph the equation with a diameter that has endpoints at (-3, 4) and (5, -2). Label the center and at least four points on the circle. Write the equation of the circle.

OpenStudy (anwaarullah):

www.mathsisfun.com/algebra/circle-equations.html

OpenStudy (zpupster):

\[(\frac{ x₂+x₁ }{ 2 } ,\frac{ y₂+y₁ }{ 2 })\] Use this to find midpoint-center

OpenStudy (zpupster):

use midpoint and an endpoint to find r and then plug into equation of circle \[(x-h)²+(y-k)² = r²\]

OpenStudy (zpupster):

graph at desmos.com